# HAIFIT: Human-Centered AI for Fashion Image Translation

### v1

@@ IV-A Dataset
Commonly available fashion datasets like Fashion FashionGEN [15] and DeepFashion [12] primarily consist of images directly collected from real-life photographs. As a consequence, these images encompass diverse backgrounds and characters, posing a challenge in constructing paired sketch-image datasets for cross-modal image translation tasks. The comparison between these datasets is illustrated in Fig. 5. To overcome this limitation, we introduce the HAIFashion dataset, comprising 3100 fashion clothing images and 933 pairs of sketch-image data meticulously hand-drawn by professional fashion designers. We employed a sketch generation algorithm to convert clothing images into corresponding sketch-image pairs, facilitating our experiments.
@@ IV-B Implementation Details
We implement our model on Pytorch and all experiments are performed on a single NVIDIA GeForce RTX4090 GPU. We used the Adam optimizer with a learning rate of 1e-4 for the generator and 5e-4 for the discriminator. The learning rates decayed at the halfway point every 100 epochs. The training process continued until the SSIM [20] showed no improvement for 10 iterations. We use the HAIFashion dataset and divide it into 2500 (600) for model training and testing respectively, and the batch size is set to 8 for all experiments.
@@ IV-C Evaluation Metrics
We employ four commonly used quantitative evaluation metrics: PSNR (Peak Signal-to-Noise Ratio) measures image reconstruction quality by comparing mean square error. SSIM (Structural Similarity Index) [20] assesses similarity between images by considering their brightness, contrast, and structure. LPIPS (Learned Perceptual Image Patch Similarity) [24] uses deep convolutional neural networks to learn perceptual differences by partitioning images into patches. FID (Fréchet Inception Distance) [4] measures dissimilarity in feature distribution and statistics between generated and real samples.
@@ IV-D Baselines
In our experiments, we compared our method with nine state-of-the-art baseline approaches. All models underwent the same preprocessing steps to ensure a fair comparison. BicycleGAN [25] focuses on modeling a distribution of possible outputs in a conditional generative model. pix2pix [7] utilizes conditional GAN to convert input images between different domains. MUNIT [6] decomposes image representations into content and style properties. UAGITI [1] incorporates attention networks to minimize the divergence of the data-generating distribution. pix2pixHD [19] utilizes a novel adversarial loss and new multi-scale architectures. USTP [11] tackles translation through shape translation and content enrichment stages. UGATIT [9] employs an attention module and a learnable normalization function. Palette [16] proposes a unified conditional diffusion model. BBDM [10] presents a novel approach utilizing the Brownian bridge diffusion model.
@@ IV-E Comparison Results
TABLE I: Overall performance on sketch-to-image generation. ($\uparrow$ indicates the higher value, the better effect. $\downarrow$ indicates the lower value, the better effect.)
Methods HAIFashion
@@
BicycleGAN [25] 12.3482 0.4258 0.1976 70.6381
pix2pix [7] 19.2761 0.3826 0.1180 48.3830
MUNIT [6] 13.3265 0.4843 0.1947 74.2733
UAGITI [1] 11.8259 0.0901 0.2383 83.8309
pix2pixHD [19] 16.1863 0.5622 0.1026 48.0041
USTP [11] 13.2456 0.5126 0.1429 58.8756
UGATIT [9] 12.8841 0.4799 0.1783 83.9190
Palette [16] 7.8216 0.0787 0.3591 116.9303
BBDM [10] 14.9788 0.5343 0.1676 59.7919
HAIFIT (Ours) 19.0660 0.6338 0.0778 28.5032
We conducted quantitative evaluations to compare both the realism and faithfulness of the generated images. The quantitative results are summarized in Table I, despite achieving the second position in the PSNR index, it is important to note that this result is primarily due to the nature of PSNR, which emphasizes the signal-to-noise ratio and is sensitive to local noise and distortion in the image. HAIFIT, on the other hand, prioritizes retaining more details from the input sketches, including the initial outline, which may lead to a slight decrease in the PSNR index. However, overall, HAIFIT demonstrates excellent performance in all four metrics. This demonstrates the effectiveness of our method in sketch-to-image generation. This success can be attributed to the integration of our proposed MFFE and pyramid generator with CSC. The MFFE helps our model learn object sketch contours and consider correlations between adjacent lines from different directions, enhancing stroke variation perception. Meanwhile, the pyramid generator with CSC can capture specific details in key regions of complex sketches, facilitating the learning of the complete structure and fine-tuning of intricate details. The results confirm that our model can well adapt to sketches with specific detailed styles and generate more realistic images.
As shown in Fig. 6, it is evident that our proposed method outperforms other baselines by delivering high-definition and diverse images. For example, although BcycleGAN, USTP, and UGATIT produce some satisfactory results, the majority of their images exhibit a single color tone and lack fine details. Similarly, pix2pix, MUNIT, and pix2pixHD demonstrate good color diversity, but suffer from issues such as uneven color distribution and blurred details. UAGITI also faces challenges related to single color and uneven color distribution. The diffusion models Palette appears severe background coloring while BBDM exhibits a uniform color distribution within the outline but there is a struggle to retain fine details in the generated images. Overall, existing methods encounter challenges in capturing specific intricate details in sketches. In contrast, our proposed method excels in handling complex sketches, leading to images with more realistic visual effects.
Figure 7: Comparison of model training and inference times.
Fig. 7 illustrates the model training and inference times for our model and the nine baselines. Our model exhibits a clear advantage in training time compared to GAN-based methods. Additionally, when compared to the methods based on the Diffusion model, our model demonstrates an inference time improvement of over 100 times. This significant time-saving advantage can greatly benefit designers by reducing costs, improving work efficiency, and increasing productivity.
IV-F Ablation study
We conduct ablation studies to analyze the effectiveness of (i) the abstract feature representation module (AFRM) and (ii) the Cross-level Skip Connection (CSC).
TABLE II: The quantitative experimental results on ablation study.
Methods HAIFashion
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
w/o AFRM and CSC 19.7086 0.6324 0.0812 30.5260
@@ IV-F Ablation study
As indicated in Table II, the removal of the CSC between different layers of HAIFIT negatively impacts our model’s performance. CSC is essential in considering feature correlations across scales and combining different receptive fields, which helps preserve important semantic details. The increase in the PSNR index after removing CSC is attributed to the loss of sketch detail features, resulting in smoother perceived details. Moreover, removing the AFRM also led to a decline in our model’s performance. AFRM learns long-term dependencies from different perspectives of the sketch using the LSTM module. It takes into account the correlation between adjacent lines in the sketch to increase the details and ensure the consistency of the generated image. Notably, when replacing the LSTM module with the Transformer [18] module, all four performance indicators of the model deteriorated. This decline can be attributed to the temporal dependence offered by the LSTM module, ensuring consistent feature extraction across all directions of the sketch and yielding a more intricate and coherent feature representation. In contrast, the Transformer module, relying on the self-attention mechanism, captures relationships within sequences but lacks strong modeling capabilities in preserving the sequence’s order of features. This outcome underscores the significance of considering the temporal correlation between lines to enhance model performance. Based on the above analysis, the AFRM and CSC modules enhance our model’s capability to handle complicated sketches. Consequently, HAIFIT enables the generation of high-quality and diverse clothing images while maintaining the design style of the sketches.
@@
The qualitative results of the ablation study on sketch-to-image generation are presented in Fig. 8. It can be observed that removing the AFRM or CSC modules leads to uneven color distribution in the generated images or loss of certain details. Moreover, when both modules are removed simultaneously, these issues become more pronounced. While replacing the Transformer module may also get good results, problems can still be observed in the image color distribution. These findings further confirm the effectiveness of our model.
@@ V Conclusions
In this paper, we introduce HAIFIT, a creative sketch-to-image generation model tailored for fashion design. We effectively enhance both the global consistency of feature vectors and the richness of intricate textures in the latent space through the proposed MFFN and CSC. This enhancement enables the generation of high-quality, realistic clothing images. Furthermore, we present for the first time a dataset of sketches drawn by professional clothing designers, and based on this, obtained the HAIFashion dataset. We conduct extensive experiments comparing our method with state-of-the-art approaches. The qualitative and quantitative results robustly demonstrate the effectiveness of our proposed methodology.

### v2

@@ IV-A HAIFashion Dataset
Considering our objective of learning clothing sketch features and generating high-quality clothing images while preserving the designer’s style and intricate details, it’s imperative to construct a new clothing dataset comprising paired images and sketches for model training and testing. While directly hand-drawing a large dataset of garment image-sketch pairs by designers may seem challenging, a more feasible approach involves designers creating a small amount of clothing image-sketch paired data, which can then be expanded using an image-to-sketch generation algorithm. Thanks to the personalized sketch generation algorithm proposed by StyleMe [31], this dataset construction becomes feasible. Thus, our dataset construction task focuses on building a fashion clothing image dataset for image-to-sketch generation.
Common publicly available fashion datasets like FashionGEN [23] and DeepFashion [19] predominantly feature realistic fashion images sourced directly from real-world photos. These images often include various backgrounds and individuals wearing the clothing in social settings. However, images with backgrounds pose significant challenges in generating single clothing sketches, which may not meet our requirements. The comparison between these datasets is illustrated in Fig. 5. In this study, we introduce a self-collected clothing dataset named HAIFashion. HAIFashion comprises 3100 pairs of fashion clothing image-sketch datasets. We achieved this by crawling 3100 background-free fashion clothing images from the internet and employing the algorithm [31] to generate designer-style clothing sketches, facilitating our experiments.
@@ IV-B Implementation Details
We implement our model on Pytorch and all experiments are performed on a single NVIDIA GeForce RTX4090 GPU. We utilize the Adam optimizer with a learning rate of 1e-4 for the generator and 5e-4 for the discriminator, the batch size is set to 8. The learning rates decayed at the halfway point every 100 epochs. The training process continued until the SSIM [30] showed no improvement for 10 iterations. For the dataset, we use our proposed HAIFashion (§ IV-A) and divide it into 2500 (600) for model training and testing respectively.
TABLE I: The Overall sketch-to-image generation performance on the HAIFashion dataset. ($\uparrow$ indicates the higher value, the better effect. $\downarrow$ indicates the lower value, the better effect.)
Methods Venue Architecture Params Structural Perceptual
(MB) PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
BicycleGAN [35] NeurIPS’17 GAN 219.2 12.3482 0.4258 0.1976 70.6381
pix2pix [12] CVPR’17 GAN 45.6 19.2761 0.3826 0.1180 48.3830
MUNIT [11] ECCV’18 GAN 120.3 13.3265 0.4843 0.1947 74.2733
UAGITI [1] NeurIPS’18 GAN 2.4 11.8259 0.0901 0.2383 83.8309
pix2pixHD [29] CVPR’18 GAN 729.8 16.1863 0.5622 0.1026 48.0041
USTP [18] ECCV’20 GAN 31.4 13.2456 0.5126 0.1429 58.8756
UGATIT [15] ICLR’20 GAN 2764.8 12.8841 0.4799 0.1783 83.9190
Palette [24] SIGGRAPH’22 DM 250.7 7.8216 0.0787 0.3591 116.9303
BBDM [17] CVPR’23 DM 1,945.6 14.9788 0.5343 0.1676 59.7919
HAIFIT (Ours) - GAN 769.6 19.0660 0.6338 0.0778 28.5032
@@ IV-C Evaluation Metrics
We employ four commonly used quantitative evaluation metrics:
  • PSNR (Peak Signal-to-Noise Ratio) assesses image quality by comparing the peak signal-to-noise ratio between the original image and the generated image. Higher PSNR values indicate smaller differences between the images and higher-quality generated images.
  • SSIM (Structural Similarity Index) [30] evaluates image quality by comparing the structural similarity between the original image and the generated image. It considers similarities in brightness, contrast, and structure to calculate the final score. Larger SSIM values signify higher similarity between the images and better generation effects.
  • LPIPS (Learned Perceptual Image Patch Similarity) [33] assesses image quality by utilizing a deep learning model and the L2 norm to calculate the perceptual difference between the generated image and the original image in the feature space. Lower LPIPS values suggest better-quality images.
  • FID (Fréchet Inception Distance) [8] evaluates image quality by comparing the statistical distribution in the feature space of the generated image with that of the original image. It employs the Fréchet distance to measure the similarity between the distributions. Lower FID values indicate higher-quality resulting images.
The first two metrics primarily gauge the completeness and structural similarity of the generated images, whereas the latter two metrics focus more on assessing the diversity, authenticity, and perceptual differences of the generated images.
@@ IV-D Baselines
In our experiments, we compared our method with nine state-of-the-art baseline approaches. All models undergo identical preprocessing steps and are trained directly following the official open-source code to ensure a fair comparison.
  • BicycleGAN [35] focuses on modeling a distribution of possible outputs in a conditional generative model, enhancing the diversity of generated images.
  • pix2pix [12] implements image translation using conditional GAN to convert input images between different domains.
  • MUNIT [11] achieves image synthesis by decomposing image representations into content and style properties.
  • UAGITI [1] incorporates attention networks to minimize divergence between relevant parts of the data-generating distribution.
  • pix2pixHD [29] utilizes a novel adversarial loss and new multi-scale generator and discriminator architectures in conditional GAN for high-resolution image synthesis.
  • USTP [18] learns from unpaired sketch and photo data and tackles synthesis through shape translation and content enrichment stages.
  • UGATIT [15] employs an attention module and a learnable normalization function to enhance the quality and diversity.
  • Palette [24] proposes a unified framework for image-to-image translation based on conditional diffusion models, demonstrating the importance of self-attention.
  • BBDM [17] presents a novel image-to-image translation approach utilizing the Brownian bridge diffusion model, which enables direct learning of domain translation through a bidirectional diffusion process.
@@ IV-E Comparison Results
Quantitative Experiments. We conducted quantitative evaluations to compare both the realism and faithfulness of the generated images. The quantitative results are summarized in Table I, despite achieving the second position in the PSNR index, it is important to note that this result is primarily due to the nature of PSNR, which emphasizes the signal-to-noise ratio and is sensitive to local noise and distortion in the image. HAIFIT, on the other hand, prioritizes retaining more details from the input sketches, including the initial outline, which may lead to a slight decrease in the PSNR index. However, overall, HAIFIT demonstrates excellent performance in all four metrics. This demonstrates the effectiveness of our method in sketch-to-image generation. This success can be attributed to the integration of our proposed MFFE and pyramid generator with CSCM. The MFFE helps our model learn object sketch contours and consider correlations between adjacent lines from different directions, enhancing stroke variation perception. Meanwhile, the pyramid generator with CSCM can capture specific details in key regions of complex sketches, facilitating the learning of the complete structure and fine-tuning of intricate details. The results confirm that our model can well adapt to sketches with specific detailed styles and generate more realistic images.
Qualitative Experiments. As shown in Fig. 6, it is evident that our proposed method outperforms other baselines by delivering high-definition and diverse images. For example, although BcycleGAN, USTP, and UGATIT produce some satisfactory results, the majority of their images exhibit a single color tone and lack fine details. Similarly, pix2pix, MUNIT, and pix2pixHD demonstrate good color diversity, but suffer from issues such as uneven color distribution and blurred details. UAGITI also faces challenges related to single color and uneven color distribution. The diffusion models Palette appears severe background coloring while BBDM exhibits a uniform color distribution within the outline but there is a struggle to retain fine details in the generated images. Overall, existing methods encounter challenges in capturing specific intricate details in sketches. In contrast, our proposed method excels in handling complex sketches, leading to images with more realistic visual effects.
Figure 7: Comparison of model training and inference times.
Time Costs. Fig. 7 illustrates the model training and inference times for our model and the nine baselines. Our model exhibits a clear advantage in training time compared to GAN-based methods. Additionally, when compared to the methods based on the Diffusion model, our model demonstrates an inference time improvement of over 100 times. This significant time-saving advantage can greatly benefit designers by reducing costs, improving work efficiency, and increasing productivity.
IV-F Ablation study
We conduct ablation studies to analyze the effectiveness of (i) the Abstract Feature Representation Module (AFRM) and (ii) the Cross-level Skip Connection Module (CSCM).
Quantitative Experiments. As indicated in Table II, the removal of the CSCM between different layers of HAIFIT negatively impacts our model’s performance. CSCM is essential in considering feature correlations across scales and combining different receptive fields, which helps preserve important semantic details. The increase in the PSNR index after removing CSCM is attributed to the loss of sketch detail features, resulting in smoother perceived details. Moreover, removing the AFRM also led to a decline in our model’s performance. AFRM learns long-term dependencies from different perspectives of the sketch using the LSTM module. It takes into account the correlation between adjacent lines in the sketch to increase the details and ensure the consistency of the generated image. Notably, when replacing the LSTM module with the Transformer [28] module, all four performance indicators of the model deteriorated. This decline can be attributed to the temporal dependence offered by the LSTM module, ensuring consistent feature extraction across all directions of the sketch and yielding a more intricate and coherent feature representation. In contrast, the Transformer module, relying on the self-attention mechanism, captures relationships within sequences but lacks strong modeling capabilities in preserving the sequence’s order of features. This outcome underscores the significance of considering the temporal correlation between lines to enhance model performance. Based on the above analysis, the AFRM and CSCM modules enhance our model’s capability to handle complicated sketches. Consequently, HAIFIT enables the generation of high-quality and diverse clothing images while maintaining the design style of the sketches.
TABLE II: The quantitative experimental results on ablation study.
Methods Structural Perceptual
@@ IV-F Ablation study
w/o AFRM and CSCM 19.7086 0.6324 0.0812 30.5260
@@
Qualitative Experiments. The qualitative results of the ablation study on sketch-to-image generation are presented in Fig. 8. It can be observed that removing the AFRM or CC modules leads to uneven color distribution in the generated images or loss of certain details. Moreover, when both modules are removed simultaneously, these issues become more pronounced. While replacing the Transformer module may also get good results, problems can still be observed in the image color distribution. These findings further confirm the effectiveness of our model.
@@ V Conclusions
In this paper, we introduce HAIFIT, a creative sketch-to-image generation model tailored for fashion design. We effectively enhance both the global consistency of sketch feature vectors and the richness of intricate textures in the latent space through the proposed MFFN and CSCM. This enhancement enables the generation of high-quality, realistic clothing images. Furthermore, We present a fashion clothing dataset HAIFashion for image translation to fill the gap in this field. We conduct extensive experiments comparing our method with state-of-the-art approaches. The qualitative and quantitative results robustly demonstrate the effectiveness of our proposed methodology in preserving sketch details and generating realistic images.